OpenStudy (anonymous):

Write an equation for the line that is parallel to the given line and passes through the given point. y = x + 8; (4, 10) A. y = x B. y = x + 6 C. y = x – 6 D. y = x + 6

OpenStudy (dan815):

this is like the most asked question on here lol

OpenStudy (ddcamp):

From the given equation, the slope (m) equals 1. Plug that into the point slope equation and solve for y \[y - y₁ = m(x-x₁)\]

OpenStudy (anonymous):

i don't understand

OpenStudy (ddcamp):

y - 10 = 1(x - 4)

OpenStudy (anonymous):

i still don't understand omg i'm sorry hahaha